# On the Bias of Next-Token Predictors Toward Systematically Inefficient Reasoning: A Shortest-Path Case Study

### v1

@@ Abstract
Recent advances in natural language processing highlight two key factors for improving reasoning in large language models (LLMs): (i) allocating more test-time compute tends to help on harder problems but often introduces redundancy in the reasoning trace, and (ii) compute is most effective when reasoning is systematic and incremental, forming structured chains of thought (CoTs) akin to human problem-solving. To study these factors in isolation, we introduce a controlled setting based on shortest-path tasks in layered graphs. We train decoder-only transformers on question–trace–answer triples using a custom tokenizer, comparing models trained on optimal bottom-up dynamic programming traces with those trained on longer, valid traces involving backtracking. Surprisingly, with the same training-token budget, models trained on inefficient traces generalize better to unseen graphs. This benefit is not due to length alone—injecting arbitrary redundancy into reasoning traces fails to help and can even hurt performance. Instead, we find that generalization correlates with the model’s confidence in next-token prediction, suggesting that long, coherent, and locally incremental traces make the training signal easier to optimize.
@@ 1 Introduction
Modern LLMs have made remarkable strides in tasks requiring reasoning and multi-step problem solving [1, 2, 3, 4]. Increasing evidence demonstrates that the performance of these models significantly improves when their reasoning unrolls in a step-by-step fashion, following CoTs reminiscent of how humans build their internal cognitive processes [5, 6, 7, 8]. Another milestone in the rapid development of LLMs is represented by the test-time-compute paradigm [9, 10, 11], driven by the intuition that harder problems often require more computational budget—and thus, longer CoTs. These insights have been central to the development of advanced reasoning agents capable of achieving unprecedented performance on complex tasks spanning mathematical problem solving [12, 13, 14], code generation [15, 16], and scientific inquiry [17, 18]. Despite such progress, reasoning remains an elusive concept—difficult to define precisely and challenging to study in the wild. How to teach machines to reason effectively, generalize across domains, and adapt to novel problems continues to be a fundamental open question.
Motivated by the need for a well-defined and interpretable setting to characterize reasoning in transformers, and inspired by [19], we turn to a controlled algorithmic task (see Fig 1): a synthetic shortest-path problem on layered graphs. Each problem instance—posed as a question to a transformer—consists of a source-to-target graph with integer edge costs; the model is tasked with answering with any minimum-cost path. While a bottom-up dynamic programming (DP) approach is the canonical optimal solution for this problem, several alternative strategies can also reach the correct answer. We design a family of such strategies and train transformers to follow them. This playground allows us to (i) generate possibly unlimited problem instances of varying levels of difficulty; (ii) probe key properties of effective reasoning agents identified in the literature. In particular, through this framework, we investigate the following questions:
@@
    Thinking step-by-step: We find that training transformers to produce intermediate steps between question (problem instance) and answer (optimal path) substantially improves performance, in line with the behavior of modern LLM on problem-solving tasks.
@@
    Next-token predictors favor inefficient traces: We motivate our findings by measuring the confidence of trained models on the prediction of the next token given the context: we find that this metric is higher for models trained on longer but systematic and locally incremental reasoning traces and lower on globally optimal strategies.
@@ 2 Problem Setup
As a testbed for our reasoning experiments, we consider a synthetic shortest-path problem in layered directed acyclic graphs (DAGs) with integer edge costs. We generate random graph instances from a family with parameters $\{L,K,C,p_{e}\}$, respectively representing the maximum number of layers, the maximum number of nodes per internal layer, the maximum edge cost, and the average connectivity between nodes in successive layers. For simplicity, the first and last layers contain exactly one node, the source and the destination of the sought path. We also ensure that no nodes are either completely disconnected from the previous layer or have no connections to the next one. Once the graph has been defined, we label the nodes in a top-to-bottom and left-to-right order. See Fig. 1 for an example of a small problem instance.
@@
The above-described task can be framed as a simple multi-step reasoning task, involving the successive solution of a set of sub-problems, i.e., the shortest path from the source node to all intermediate nodes. The goal of the reasoner is to adopt an efficient strategy for completing all the required reasoning steps and building an optimal solution. Enumeration of all partial paths would entail a $\mathcal{O}\left(K^{L}\right)$ computational cost, yet the cost can be cut down to $\mathcal{O}(LK^{2})$ if the sub-problems are conveniently ordered and the partial solutions are stored away. A bottom-up dynamic programming (DP) approach, solving the sub-problems in the layer order– from shorter to longer partial paths, yields an optimal shortest path with the minimum number of reasoning steps.
@@ 3 Experimental Setup
We leverage the controlled nature of our problem setting to define a random generator of reasoning traces with a controllable degree of efficiency. These exact traces are then used to build a training corpus of question-trace-answer triples, and train a next-token prediction model on the shortest-path task.
The parametric control over the structural efficiency of the reasoning trace is achieved through a temperature parameter $\eta$, determining the probability of prioritizing shorter vs. longer partial-path explorations, i.e., the bias of the underlying algorithm toward a layer-by-layer (positive $\eta$) or a depth-first (negative $\eta$) approach. Independent of $\eta$, when multiple partial paths of the same length can be continued, the relative exploration order is randomized. The internal logic of the exploration algorithm and its dependence on $\eta$ can be seen in Fig. 2(a). The temperature $\eta$ indirectly affects the number of reasoning steps, since previously explored paths will need revising if a better route to an intermediate node is encountered. Note, however, that the trace optimality criteria described in section 2.1 guarantee that the number of steps remains polynomial (worst-case $\mathcal{O}(CL^{2}K^{2})$  [^1] Given the structure of the exploration algorithm Fig. 2(a), each one of the $\mathcal{O}(LK)$ nodes can be at most revisited $\mathcal{O}(CL)$ times, since a best cost improvement is needed to trigger backtracking.), as evidenced from the table in Fig. 2(b).
Furthermore, we can inject reasoning redundancy, by artificially increasing the length of the trace via repetition of full reasoning steps. To study the importance of preserving the CoT structure, we consider two variants: a randomized version, where completed reasoning steps can be disregarded with a given probability, requiring (possibly multiple) additional explorations later in the trace; a systematic version, where each reasoning step is immediately repeated (possibly multiple times) and then never revisited.
@@
  • $\eta=+5$: at this temperature, the reasoning trace corresponds to a bottom-up DP trace, systematically exploring the graph in a layer-by-layer order.
@@
  • $\eta=-5$: the reasoning trace systematically prioritizes a depth-first approach, requiring substantial backtracking.
  • $\eta=+5$ efficiency, randomized redundancy (RR): in this setting, each reasoning step is ignored (and later repeated) with probability $1/2$.
  • $\eta=+5$ efficiency, deterministic redundancy (DR): in this setting, each reasoning step is systematically repeated twice, consecutively.
Note that, given the exponential law employed to sample the layer order, the temperature values $\eta=\pm 5$ are large enough to fully order the exploration protocol (effectively behaving as $\eta=\pm\infty$, but avoiding numerical instabilities). As shown in Fig. 2(b), the average length of the traces increases with lower values of the temperature $\eta$, changing by roughly $75\%$ when switching between $\eta=\pm 5$. In the redundancy cases, the number of repetitions is matched between the randomized and deterministic versions. Examples of the different trace types are provided in the Supplementary Materials (SM).
@@
3:for $n$in layer1 do
4:     $\mathcal{E}$, $(n0,n,0)$); append($\mathcal{W}$, $1$)
@@
8:     $c$, path $\leftarrow\mathcal{C}$(src) $+$ cost(src,dst), $\mathcal{P}$(src) $\cup$ dst
@@
12:              e, w $\leftarrow$ (dst,n,layer$+1$), $\exp(-\boldsymbol{\eta}($ layer+1))+1))+ 1 ) )
@@
14:                  $\mathcal{E}$, e); append($\mathcal{W}$, w)
@@
Figure 2: Dependency on the efficiency temperature $\boldsymbol{\eta}$ (a) of the exploration algorithm (line 12); (b) of the mean and standard deviation of the number of reasoning steps for the different types of traces (estimated on $100$K independent samples).
@@
We train from scratch a Phi3 [20] small language model, with 3 layers, 12 attention heads, 768 hidden dimensions, and $28.5$M total parameters, for the next-token prediction task on the procedurally generated training examples. During training, we mask out the question (i.e., the graph information) and the PAD tokens from the loss function, so that the model only learns to predict traces and answers within the context of the question. We employ a constant learning rate of $2\times 10^{-5}$, and a batch size of $\sim 16$K tokens (excluding the PAD tokens for a fair comparison between efficiency levels). In SM, we show how adding layers to the transformer architecture can impact performance and sample efficiency.
@@
We also check for multiple secondary metrics on the quality of the reasoning steps in the trace, including: i) if they contain syntax errors; ii) if they are incremental; iii) if they only build upon optimal partial paths; iv) if the path-cost declaration is consistent; iv) if the reasoning steps are repeated. To simplify the exposition, the analysis of these metrics is deferred to SM.
Finally, inspired by [21], we also measure the next-token confidence of the model along the reasoning trace and answer, i.e., the average sampled token probability. Note that here we find this metric to be equivalent to the min-margin metric suggested in [21], in agreement with a footnote observation therein.
@@ 4 Results
Figure 3: Learning to find the shortest path. (a) Generalization performance of two models trained on $\sim 340$K graphs, respectively without reasoning traces (dashed) and with the optimal $\eta=+5$ traces (full), over graphs with depths $3$-$5$-$7$.   (b) Progress on intermediate training goals for the $\eta=+5$ model.   (c) Acquisition of the integer addition sub-task, during the $\eta=+5$ training. The plot shows the probability of the model predicting the correct $row+column$ sums at different epochs.
The first finding we report on, shown in Fig. 3(a), is that a next-token predictor can learn to solve shortest-path problems in moderate-sized graph instances, when sufficient training data is presented (here $\sim 340$K graph examples). However, good generalization on the larger graph instances can only be achieved when the model is allowed to produce a reasoning trace before constructing an optimal path: the solid curves show the performance, on $3,5,7$ layer graphs, of a model trained in the dynamic programming limit $\eta=+5$. Instead, the dashed curves show that a model trained on the question-answer task, with the same number of examples, is unable to learn a generalizable rule for solving previously unseen large problems.
In Fig. 3(b), we break down the learning process for the $\eta=+5$ model, showing the pace at which intermediate learning goals are unlocked, such as proposing candidate solutions that are possible, or correctly associating paths and costs. In Fig. 3(c), we study how the model acquires the sub-task of integer addition (here restricted to a small subset $\mathcal{O}(CL)$ of possible cumulative sums). In the figure, we measure the probability of the model assigning the maximum logit to the correct sum token, and display the order in which the different additions are learned during training. Note that this subtask is not trivial [22, 23], although the training set extensively covers the relevant range of cumulative costs (see SM), since the model has to understand where to pick up the partial costs to be added from the question and the previous steps in the trace.
@@
Figure 4: Impact of trace efficiency. (a) Comparison of the generalization performance between models trained on efficient $\eta=+5$, intermediate $\eta=0$, and inefficient $\eta=-5$ traces, with a training token budget of $32$M (dashed) and $128$M (full) tokens.   (b) Next-token confidence measured on the same training trajectories of models trained on efficient $\eta=+5$, intermediate $\eta=0$, and inefficient $\eta=-5$ traces with a training token budget of $128$M.   (c) Training losses for 3 different seeds of $\eta=-5$, showing sudden jumps at the beginning of the 2nd epoch, and of $\eta=+5$, where the optimization is slower and more continuous.
@@
We explore the impact of the reasoning trace efficiency on model performance, at fixed token budget (dashed $32$M tokens, full $128$M tokens) during training. In Fig. 4(a), the best accuracy in predicting the shortest path is achieved by the $\eta=-5$ model (blue), with the longest reasoning traces and systematic backtracking, followed by the DP-like $\eta=+5$ (red), and finally by the $\eta=0$ models (green). Note that, since the token budget at training is fixed, the $\eta=-5$ inefficient model sees about $1/3$ less graph examples compared to the efficient one at $\eta=+5$, yet absorbs the training set information more effectively. Moreover, notice that more test-time compute –here the length of the trace– does not necessarily imply better test performance, as can be seen from the sub-optimal $\eta=0$ curves. In Fig. 1(c), we also train the $\eta=\pm 5$ models with an equal number of graph examples ($\sim 200$K, as in the $128$M dataset for $\eta=-5$), finding an even larger performance gap.
In Fig. 4(b), we find a plausible explanation of the effectiveness of $\eta=-5$ and the poor performance of $\eta=0$, looking at the next-token confidence [21] of the three models. While also the $\eta=0$ traces are longer than the $\eta=+5$ ones, the associated flat distribution over the exploration order, mixing depth-first and layer-by-layer exploration, is harder to understand and mimic for the trained model. This can be seen from the low confidence on the next-token prediction task of the $\eta=0$ model. On the other hand, both $\eta=+5$ and $\eta=-5$ strategies share a deterministic approach in the layer exploration order. However, while the DP-like algorithm is globally more efficient and logical, the depth-first approach entails longer trains of thought, with incremental steps that immediately build upon previous steps, resulting in a more consequential reasoning process. This “inertia” in the exploration helps the model predict what to do next, with an increase in next-token confidence and a beneficial effect on training outcome and performance. A similar plot, but at fixed number of training graph examples, is shown in Fig. 1(d).
@@
The effect of trace predictability can also be seen from the optimization trajectories, in Fig. 4(c). The $\eta=-5$ trajectories often show a sudden jump in the training loss, around the beginning of epoch 2, highlighting an “eureka” transition in the interpretation of the reasoning trace examples that unlocks better generalization. During the jump, the model transitions from repeating the same reasoning steps from $20\%$ to $3\%$ of the times (note that, in $40\%$ of the cases, repeated cost computations lead to different costs), considering optimal partial paths from $50\%$ to $98\%$ of the steps, and producing costs consistent with the paths from $33\%$ to $97\%$ of the steps. We hypothesize this sudden transition could be explained by the emergence of specific circuits within the model forward pass, similar to those identified by [24, 25, 26, 27, 28]. We see the mechanistic interpretability analysis of our trained model as an interesting avenue for future work.
@@
To further explore the impact of increased test-time compute, in the absence of stochastic confounders that decrease the predictability of the trace, we compare the $\eta=+5$ baseline with a model trained on deterministically redundant $\eta=+5$ DR traces. Since we repeat each reasoning step twice in a row, in principle, the model can build a mechanism for choosing the next path that relies on this repetition and increases its computational budget, developing the computation horizontally –in the sequence length dimension. Note that the elongated CoTs have more steps than $\eta=-5$ on average. In Fig. 5(a), we show that artificially increasing the length of the reasoning traces, without a systematic change in the exploration strategy, induces a slight performance deterioration if trained with a fixed token budget ($128$M). This aligns with recent research showing that CoT length is task-dependent [29] and that the global structure of the CoT is often more important than its content [30].
@@
0.0 18 $231\pm 136$
0.2 39 $175\pm 121$
0.5 76 $94\pm 70$
0.7 80 $82\pm 56$
1.0 81 $74\pm 48$
1.2 75 $72\pm 46$
1.5 60 $71\pm 44$
1.7 44 $68\pm 43$
2.0 19 $65\pm 42$
Figure 5: Redundant traces. (a) Comparison of generalization performance between models trained on traces with efficiency $\eta=+5$, $\eta=+5$DR, and $\eta=+5$RR (with sampling temperatures $T=1$ (dashed) and $T=0$ (full)), trained with a $128$M token budget.   (b) Regularization effect of sampling temperature on $\eta=+5$RR, where the answer accuracy improves and the average CoT length converges to the expected one from training data at higher temperatures.
@@
In the previous experiments, a non-systematic structure in the reasoning trace—as in the case of $\eta=0$—was associated with a reduced capability of the model to predict the next token, affecting both optimization and generalization performance. Structural perturbations to the reasoning traces can, in fact, strongly hinder performance [31]. A similar effect can be seen if the redundancy is introduced in a randomized fashion. In Fig. 5(a), we show the probability of producing a correct shortest path solution for a model trained in the $\eta=+5$ RR setting and generating at zero temperature. Apart from the reduction in the model accuracy compared to the deterministic analogous, in Fig. 5(b) we also observe that the generated trace length diverges from the average one in the training set: the model seems to get stuck in repetition loops that can elongate the CoT indefinitely.
A similar cause can be traced for the poor performance in the $\eta=0$ case, in Fig. 6(a). At zero sampling temperature, the model seems to gravitate towards the trace length of $\eta=-5$, showing a preference for long CoTs, although the training examples also propose shorter traces with a majority of layer-by-layer steps. In both cases, the models display a bias towards mechanisms that systematically induce longer, yet more autocorrelated, traces. This finding can be connected with multiple works showing the bias of LLMs for high verbosity [2, 32, 33].

### v2

@@ Abstract
Recent advances in natural language processing highlight two key factors for improving reasoning in large language models (LLMs): (i) allocating more test-time compute tends to help on harder problems but often introduces redundancy in the reasoning trace, and (ii) compute is most effective when reasoning is systematic and incremental, forming structured chains of thought (CoTs) akin to human problem-solving. To study these factors in isolation, we introduce a controlled setting based on shortest-path tasks in layered graphs. We train decoder-only transformers on question–trace–answer triples using a custom tokenizer, comparing models trained on optimal bottom-up dynamic programming traces with those trained on longer, valid traces involving backtracking. Surprisingly, with the same training-token budget, models trained on inefficient traces generalize better to unseen graphs. This benefit is not due to length alone—injecting arbitrary redundancy into reasoning traces fails to help and can even hurt performance. Instead, we find that generalization correlates with the model’s confidence in next-token prediction, suggesting that long, coherent, and locally incremental traces make the training signal easier to optimize. The code is available at https://github.com/riccardoalberghi/DP
@@ 1 Introduction
Modern LLMs have made remarkable strides in tasks requiring reasoning and multi-step problem solving openai2024openaio1card ; deepseekai2025deepseekr1incentivizingreasoningcapability ; qwq32b ; skywork-or1-2025 . Increasing evidence demonstrates that the performance of these models significantly improves when their reasoning unrolls in a step-by-step fashion, following CoTs reminiscent of how humans build their internal cognitive processes wei2022chain ; wang2022selfconsistency ; yeo2025demystifyinglongchainofthoughtreasoning ; li2024chainthoughtempowerstransformers . Another milestone in the rapid development of LLMs is represented by the test-time-compute paradigm snell2024scalingllmtesttimecompute ; muennighoff2025s1 ; hou2025advancinglanguagemodelreasoning , driven by the intuition that harder problems often require more computational budget—and thus, longer CoTs. These insights have been central to the development of advanced reasoning agents capable of achieving unprecedented performance on complex tasks spanning mathematical problem solving openai2024gpt4ocard ; yang2024qwen25mathtechnicalreportmathematical ; shao2024deepseekmathpushinglimitsmathematical , code generation jiang2024surveylargelanguagemodels ; hui2024qwen2 , and scientific inquiry lu2024aiscientist ; majumder2024discoverybenchdatadrivendiscoverylarge . Despite such progress, reasoning remains an elusive concept—difficult to define precisely and challenging to study in the wild. How to teach machines to reason effectively, generalize across domains, and adapt to novel problems continues to be a fundamental open question.
Motivated by the need for a well-defined and interpretable setting to characterize reasoning in transformers, and inspired by lehnert2024beyondastar , we turn to a controlled algorithmic task (see Fig 1): a synthetic shortest-path problem on layered graphs. Each problem instance—posed as a question to a transformer—consists of a source-to-target graph with integer edge costs; the model is tasked with answering with any minimum-cost path. While a bottom-up dynamic programming (DP) approach is the canonical optimal solution for this problem, several alternative strategies can also reach the correct answer. We design a family of such strategies and train transformers to follow them. This playground allows us to (i) generate possibly unlimited problem instances of varying levels of difficulty; (ii) probe key properties of effective reasoning agents identified in the literature. In particular, through this framework, we investigate the following questions:
@@
    Thinking step-by-step: We find that training transformers to produce intermediate steps between question (graph instance) and answer (optimal path) substantially improves performance, in line with the behavior of modern LLM on problem-solving tasks.
@@
    Next-token predictors favor inefficient traces: We motivate our findings by measuring the confidence of trained models in predicting the next token. We find that this metric is higher for models trained on longer but systematic and locally incremental reasoning traces and lower on globally optimal strategies.
@@ 2 Problem Setup
As a testbed for our reasoning experiments, we consider a synthetic shortest-path problem in layered directed acyclic graphs (DAGs) with integer edge costs. We generate random graph instances from a family with parameters $\{L,K,C,p_{e}\}$, respectively representing: the maximum number of layers, the maximum number of nodes per internal layer, the maximum edge cost, and the average connectivity between nodes in successive layers. For simplicity, the first and last layers contain exactly one node, the source and the destination of the sought path. We also ensure that no nodes are either completely disconnected from the previous layer or have no connections to the next one. Once the graph has been defined, we label the nodes in a top-to-bottom and left-to-right order. See Fig. 1 for an example of a small problem instance.
@@
The above-described task can be framed as a simple multi-step reasoning task, involving the successive solution of a set of sub-problems, i.e., the shortest path from the source node to all intermediate nodes. The goal of the reasoner is to adopt an efficient strategy for completing all the required reasoning steps and building an optimal solution. Enumeration of all partial paths would entail a $\mathcal{O}\left(K^{L}\right)$ computational cost, yet the cost can be cut down to $\mathcal{O}(LK^{2})$ if the sub-problems are conveniently ordered and the partial solutions are stored away. A bottom-up dynamic programming (DP) approach, which solves the sub-problems in layer order—from shorter to longer partial paths, yields an optimal shortest path with the minimum number of reasoning steps.
@@ 3 Experimental Setup
We leverage the controlled nature of our problem setting to define a random generator of reasoning traces with a tunable degree of efficiency. These exact traces are then used to build a training corpus of question-trace-answer triples, and train a next-token prediction model on the shortest-path task.
The efficiency level of the CoTs is uniquely determined by the exploration order according to which the model computes partial paths and costs and then composes them to build the optimal path. As the trace generator traverses the graph, it maintains an exploration queue containing path continuations yet to be considered. The associated priority weights depend on a temperature parameter, the efficiency $\eta$, controlling whether shorter vs. longer partial paths should be explored first. When multiple partial paths of the same length can be continued, their relative order is fully randomized. The internal logic of the exploration algorithm and its dependence on $\eta$ can be seen in Fig. 2(a).
Thus, $\eta$ can bias the underlying algorithm toward a layer-by-layer (positive $\eta$) or a depth-first (negative $\eta$) approach. This directly affects the number of reasoning steps, since previously explored paths will need revising if a better route to an intermediate node is encountered. For this reason in the paper we refer to positive $\eta$ values as efficient (less total steps) and to negative $\eta$ values as inefficient (more total steps). Note, however, that the trace optimality criteria described in section 2.1 guarantee that the computational cost number of steps remains polynomial (worst-case $\mathcal{O}(CL^{2}K^{2})$  [^1] Given the structure of the exploration algorithm Fig. 2(a), each one of the $\mathcal{O}(LK)$ nodes can be at most revisited $\mathcal{O}(CL)$ times, since a best cost improvement is needed to trigger backtracking.), as evidenced from the table in Fig. 2(b).
Furthermore, we can inject reasoning redundancy, by artificially increasing the length of the trace via repetition of full reasoning steps. To study the importance of preserving the CoT structure, we consider two variants: a deterministic version, where each reasoning step is immediately repeated and then never revisited and a randomized version, where completed reasoning steps are re-appended to the exploration queue with probability 1/2.
@@
  • $\eta=+5$ (DP): at this temperature, the reasoning trace corresponds to a bottom-up DP trace, systematically exploring the graph in a layer-by-layer order.
@@
  • $\eta=-5$ (DFS): the reasoning trace systematically prioritizes a depth-first approach, requiring substantial backtracking.
  • $\eta=+5$ (DR): each reasoning steps is deterministically repeated twice in a row.
  • $\eta=+5$ (RR): each reasoning step, after completion, is re-added to the exploration queue with probability $1/2$. Note that this implies that, in expectation, each reasoning path is repeated twice.
Note that, given the exponential law employed to sample the layer order, the efficiency values $\eta=\pm 5$ are large enough to fully order the exploration protocol (effectively behaving as $\eta=\pm\infty$, but avoiding numerical instabilities). As shown in Fig. 2(b), the average length of the traces increases with lower values of the efficiency $\eta$, changing by roughly $75\%$ when switching from $\eta=+5$ (DP) to $\eta=-5$ (DFS). In the redundancy cases, the number of repetitions is matched between the randomized and deterministic versions. Examples of the different trace types are provided in the Supplementary Materials (SM).
@@
3:for $n$ $\in$ layer1 do
4:  append($\mathcal{E}$, $(n0,n,0)$); append($\mathcal{W}$, 11)
@@
8:  $c$, path $\leftarrow\mathcal{C}$(src) ++ cost(src,dst), $\mathcal{P}$(src) $\cup$ dst
@@
12:    e, w $\leftarrow$ (dst,n,layer+1+1), $\exp(-\boldsymbol{\eta}($ layer+1))+1))
@@
14:      append($\mathcal{E}$, e); append($\mathcal{W}$, w)
@@
Figure 2: Impact of the efficiency $\boldsymbol{\eta}$. (a) The exploration algorithm used for determining the shortest-path. The exploration order depends on the parameter $\eta$ (line 12). (b) The effect of $\eta$ on the distribution of number of reasoning steps (estimated on 100K independent samples).
@@
We train from scratch a Phi3 abdin2024phi3technicalreporthighly small language model, with 3 layers, 12 attention heads, 768 hidden dimensions, and 28.528.5M total parameters, for the next-token prediction task on the procedurally generated training examples. During training, we mask out the question (i.e., the graph information) and the PAD tokens from the loss function, so that the model only learns to predict traces and answers within the context of the question. We employ a constant learning rate of $2\times 10^{-5}$, and a batch size of $\sim 16$K tokens (excluding the PAD tokens for a fair comparison between efficiency levels). In SM, we show how adding layers to the transformer architecture can impact performance and sample efficiency.
@@
We also check for multiple secondary metrics on the quality of the reasoning steps in the trace, including: i) if they contain syntax errors; ii) if they are incremental; iii) if they only build upon optimal partial paths; iv) if the path-cost declaration is consistent; v) if the reasoning steps are repeated. To simplify the exposition, the analysis of these metrics is deferred to SM.
Finally, inspired by wang2024chain , we also measure the next-token confidence of the model along the reasoning trace and answer, i.e., the average sampled token probability. Note that here we find this metric to be equivalent to the min-margin metric suggested in wang2024chain , in agreement with a footnote observation therein.
@@ 4 Results
Figure 3: Learning to find the shortest path. (a) Generalization performance of two models trained on $\sim 340$K graphs, respectively without reasoning traces (dashed) and with the $\eta=+5$ (DP) traces (full), over graphs with depths 33-55-77.   (b) Progress on intermediate training goals for the $\eta=+5$ (DP) model.   (c) Acquisition of the integer addition sub-task, during the $\eta=+5$ (DP) training. The plot shows the probability of the model predicting the correct $row+column$ sums at different epochs.
The first finding we report on, shown in Fig. 3(a), is that a next-token predictor can learn to solve shortest-path problems in moderate-sized graph instances, when sufficient training data is presented (here $\sim 340$K graph examples). However, good generalization on the larger in-distribution graph instances can only be achieved when the model is allowed to produce a reasoning trace before returning an optimal path: the solid curves show the performance, on $3,5,7$ layer graphs, of a model trained in the dynamic programming limit $\eta=+5$ (DP). Instead, the dashed curves show that a model trained on the question-answer task, with the same number of examples, is unable to learn a generalizable rule for solving previously unseen large problems.
In Fig. 3(b), we break down the learning process for the $\eta=+5$ (DP) model, showing the pace at which intermediate learning goals are unlocked, such as proposing candidate solutions that are possible, or correctly associating paths and costs. In Fig. 3(c), we study how the model acquires the sub-task of integer addition (here restricted to a small subset $\mathcal{O}(CL)$ of possible cumulative sums). In the figure, we measure the probability of the model assigning the maximum logit to the correct sum token, and display the order in which the different additions are learned during training. Note that this subtask is not trivial quirke2024understandingadditiontransformers ; nanda2023progressmeasuresgrokkingmechanistic , although the training set extensively covers the relevant range of cumulative costs (see SM), since the model has to understand where to pick up the partial costs to be added from the question and the previous steps in the trace.
@@
Figure 4: Impact of trace efficiency. (a) Comparison of the generalization performance between models trained on efficient $\eta=+5$ (DP), intermediate $\eta=0$, and inefficient $\eta=-5$ (DFS) traces, with a training token budget of 3232M (dashed) and 128128M (full) tokens.   (b) Next-token confidence measured on the test set of models trained on efficient $\eta=+5$ (DP), intermediate $\eta=0$, and inefficient $\eta=-5$ (DFS) traces with a training token budget of 128128M.   (c) Training losses for 5 different seeds of $\eta=-5$ (DFS), showing sudden jumps at the beginning of the 2nd epoch, and of $\eta=+5$ (DP), where optimization is slower and more continuous.
@@
We explore the impact of the reasoning trace efficiency on model performance, at fixed token budget by comparing 3232M tokens with 128128M tokens. In Fig. 4(a), the best accuracy in predicting the shortest path is achieved by the $\eta=-5$ (DFS) model, with the longest reasoning traces and systematic backtracking, followed by the DP-like $\eta=+5$ (DP), and finally by the $\eta=0$ models. Note that, since the token budget at training is fixed, the $\eta=-5$ (DFS) inefficient model sees about $1/3$ fewer graph examples compared to the efficient one at $\eta=+5$ (DP), yet absorbs the training set information more effectively. Moreover, as shown in Fig. 2(b), the traces for $\eta=0$ are longer than those of $\eta=+5$, yet the corresponding curves are sub-optimal—highlighting that higher test-time compute, represented by longer traces, does not necessarily translate into better performance. In Fig. 1(c), we also train the $\eta=\pm 5$ models with an equal number of graph examples ($\sim 200$K, as in the 128128M dataset for $\eta=-5$), finding an even larger performance gap.
In Fig. 4(b), we find a plausible explanation of the effectiveness of $\eta=-5$ and the poor performance of $\eta=0$, looking at the next-token confidence wang2024chain of the three models. While the $\eta=0$ traces are longer than the $\eta=+5$ (DP) ones, the associated flat distribution over the exploration order, mixing depth-first and layer-by-layer exploration, reduces the confidence of the model in predicting the next step. This, in turn, undermines the learning effectiveness of the model trained on this trace type.
We can precisely quantify the degeneracy of the exploration order for each value of $\eta$, by computing the average surprisal $\mathcal{S}$ (i.e., the Shannon information) associated with the selection of the next path from the exploration queue. We obtain
$\mathcal{S}_{\eta=+5}=1.3262\pm 0.0006,\qquad\mathcal{S}_{\eta=-5}=0.4821\pm 0.0002,\qquad\mathcal{S}_{\eta=0}=1.905\pm 0.003,$
confirming that for $\eta=0$ the order of the reasoning steps is the most uncertain. On the other hand, both $\eta=+5$ (DP) and $\eta=-5$ (DFS) strategies share a deterministic approach in the layer exploration order, but the degeneracy of equivalent choices is higher for $\eta=+5$ (DP). A similar study of the gap in next-token confidence, but for a fixed number of training graph examples, is shown in Fig. 1(d).
@@
The effect of trace predictability can also be seen from the optimization trajectories, in Fig. 4(c). The $\eta=-5$ (DFS) trajectories often exhibit a sudden jump in the training loss, occurring around the beginning of epoch 2, which highlights a “eureka” transition in the interpretation of the reasoning trace examples. We hypothesize this sudden transition could be explained by the emergence of specific circuits within the model forward pass, similar to those identified by olsson2022incontextlearninginductionheads ; minegishi2025incontext ; cabannes2024iterationheadmechanisticstudy ; behrens2024understanding ; garnierbrun2024transformerslearnstructureddata .We therefore see the mechanistic interpretability analysis of our trained model as an interesting avenue for future work.
@@
To further explore the impact of increased test-time compute, in the absence of stochastic confounders that decrease the predictability of the trace, we compare the $\eta=+5$ (DP) baseline with a model trained on deterministically redundant $\eta=+5$ (DR) traces. Since we repeat each reasoning step twice in a row, in principle, the model can build a mechanism for choosing the next path that relies on this repetition for artificially increasing the test-time compute. Note that the elongated CoTs have more steps than $\eta=-5$ (DFS) on average. In Fig. 5(a), we show that artificially increasing the length of the reasoning traces, without a systematic change in the exploration strategy, induces a slight performance deterioration if trained with a fixed token budget (128128M). This aligns with recent research showing that CoT length is task-dependent wu2025lessunderstandingchainofthoughtlength and that the global structure of the CoT is often more important than its content li2025llmseasilylearnreason .
@@
0.0 8 $244\pm 182$
0.2 37 $177\pm 132$
0.5 81 $89\pm 63$
0.7 85 $77\pm 51$
1.0 84 $72\pm 44$
1.2 79 $71\pm 43$
1.5 62 $68\pm 42$
1.7 44 $67\pm 43$
2.0 20 $62\pm 43$
Figure 5: Redundant traces. (a) Comparison of generalization performance between models trained on traces with efficiency $\eta=+5$ (DP), $\eta=+5$ (DR), and $\eta=+5$ ($R$) (with sampling temperatures $T=1$ (dashed) and $T=0$ (full)), trained with a 128128M token budget.   (b) Regularization effect of sampling temperature on $\eta=+5$ ($R$), where the answer accuracy improves and the average CoT length converges to the expected one from training data at higher temperatures.
@@
In the previous experiments, a non-systematic structure in the reasoning trace—as in the case of $\eta=0$—was associated with a reduced capability of the model to predict the next token, affecting both optimization and generalization performance. Structural perturbations to the reasoning traces can, in fact, strongly hinder performance li2025llms . A similar effect can be seen if the redundancy is introduced in a randomized fashion. In Fig. 5(a), we show the probability of producing a correct shortest path solution for a model trained in the $\eta=+5$ (RR) setting, at zero sampling temperature. Apart from the reduction in the model accuracy compared to the deterministic analogue, in Fig. 5(b) we also observe that the generated trace length initially diverges from the expected one: the model enters repetition loops that can elongate the CoT indefinitely. Longer training is required for regularizing this behavior and eventually attaining better accuracy.
A similar high-verbosity tendency can be traced in the $\eta=0$ case, in Fig. 6(a). At zero sampling temperature, the model initially gravitates towards the trace lengths of $\eta=-5$ (DFS), gradually recovering after many epochs. In both cases, the models display a bias towards mechanisms that systematically induce longer traces. This finding is consistent with multiple works showing the bias of LLMs for high verbosity deepseekai2025deepseekr1incentivizingreasoningcapability ; nayab2025concisethoughtsimpactoutput ; chen2025think23overthinkingo1like .